# Cosmography with next-generation gravitational wave detectors

### v1

@@ 4.2 Bright sirens
As seen in section 2.1, BNS mergers can lead to a variety of EM counterparts which can be detected by various EM observatories (see section 3.2). In this section, we focus on kilonovae that are detected by LSST for bright siren measurement. From the simulated population of detected BNS events, we choose the sub-population that lies within the redshift of $0.5$ and can be localized in the sky to $\Delta\Omega\leq 100\,\mathrm{deg}^{2}$, corresponding to each detector network. The choice of redshift corresponds to our expectation of the range of LSST for kilonova measurement, while the $\Delta\Omega$ cut-off ensures that LSST can observe the kilonova within $10$ sky patches. Then, using the kilonova model discussed in Refs. [212, 213, 214], we generate kilonova light curves in the $ugrizy$ bands of LSST, for all the events in these sub-populations. Assuming single-exposure with an exposure time of $30$ seconds, we claim that a kilonova is detected if its peak brightness exceeds the limiting magnitude of the telescope in that particular band. Using this approach, we find that most kilonova observations correspond to the $g-$band, with the maximum redshift at which a kilonova is detected to be $z=0.28$ (validating our $z\leq 0.5$ cut-off for the sub-populations). Further, we assume that only $40\%$ of these kilonovae will actually be detected [215]. This fraction aims to account for the LSST sky coverage, systematic biases as well as duty-cycles corresponding to the GW and EM observatories. The remaining set of events, ones that are detected both in the GW and the EM band, are used for $H_{0}$ measurement using the same methodology that was used for golden dark sirens in section 4.1.
@@
We note that our bright siren estimates could be on the conservative side, as a result of our choice of the NS EOS, LSST detection strategy, or restricting ourselves to kilonova follow-up. Several studies have underscored the significance of short gamma-ray burst detections in enhancing the precision of bright siren measurements [215, 216, 149, 217]. These studies particularly emphasize their efficacy in probing events at high redshifts, thereby contributing to constraints on $\Omega_{m}$. Moreover, while our focus has been on kilonovae observations BNS mergers, a valuable avenue lies in considering NSBH mergers, which can also produce kilonova counterparts, thereby extending the pool of potential bright sirens for $H_{0}$ measurement [110, 102]. Nevertheless, amidst these considerations, the overarching conclusion is that BNS bright sirens can, by themselves, achieve $H_{0}$ measurements with sub-percent accuracy. This underscores the potential of BNS bright sirens as a powerful tool for precise cosmological parameter estimation.
@@ 4.3 High-$z$ spectral sirens
We forecast the capabilities of spectral sirens performing a hierarchical Bayesian analysis on simulated data. We follow our fiducial population of BBHs and BNSs described in the appendices and simulate their detected posterior samples following the method described in [131]. We assume a flat $\Lambda$CDM cosmology whose expansion rate is described by $H_{0}$ and $\Omega_{m}$, cf equation (1) and 2, taking as fiducial values the ones reported in Planck 2018 [199]: $H_{0}=67.66$km/s/Mpc and $\Omega_{m}=0.31$. We take a 40km CE sensitivity as representative for XG detectors. The inference is performed using numpyro [218] and accelerated with JAX [219]. Our code will be available on github. The inference is performed across all the cosmological and mass function parameters. We do not consider NSBHs as their mass spectrum is still largely unknown.
In figure 6 we present the projected constraints on the expansion rate of the Universe as a function of redshift. We compare the $1\sigma$ relative errors for a fixed number of BBHs and BNSs, taking the error with respect to the mean of the $H(z)$ traces at each redshift. We choose 50,000 events which corresponds approximately to the detection rate of BBHs per year with XG detectors (see figure 3). Since we assumed the same merger rate history $\mathcal{R}(z)$ for all compact binaries (see details in B.2) and BNS selection effects brings them to lower redshifts, the population of BNSs has more constraining power at lower $z$. Within the population of BBHs, it is the peak at low masses that dominates the inference [131]. Both BNSs and BBHs can constrain the expansion rate with sub-percent precision over a wide range of redshifts, with the best constrained values at around $z\sim 1$. A combined analysis of all CBCs will only improve these constraints and allow to more efficiently break the possible degeneracies with the redshift evolution of the mass spectrum [131].
@@ 4.4 All-in XG cosmography
Although we have introduced each type of standard siren separately for pedagogical reasons, they will all be analyzed simultaneously with XG detectors. In fact, as already discussed in the context of current detectors [115], the population inference of the spectral siren method must be intertwined with the dark siren method, since a wrong population model assumption can bias the dark siren cosmology. Recently, there have been developments of algorithms to jointly infer the GW population and cosmology with additional galaxy catalog information using hierarchical Bayesian analyses [220, 221] and neural posterior estimation [222].
@@
Figure 7: Joint inference of spectral sirens and (golden) dark sirens with one year of observation of XG detectors. On the left, posterior distribution for $h_{0}$ and $\Omega_{m}$. On the right, $1\sigma$ relative errors in $H(z)$. The dark siren prior in $H_{0}$ has a $1\%$ error and we use BBHs as the reference population for spectral sirens.
@@ 4.5 Stress-testing the standard cosmological model
For the second class, many (exotic) variants are possible. One that can particularly be well tested with GWs is that general relativity (GR) does not hold at cosmological scales. Among other things, this changes the propagation of GWs. Since compact binary coalescence occur at cosmological distances, even very small deviations in the GW propagation with respect to GR can be strongly constrained. The paradigmatic modifications of gravity at cosmological scales that have been explored are those that change the speed of propagation and the scaling of the luminosity distance with redshift [225]. In particular, for the case of scalar-tensor theories, the tremendous power of tests of the speed of gravity were anticipated [226, 227] and spectacularly demonstrated with the first multi-messenger event, GW170817 [228, 229, 230, 231]. After the tight constraints on the speed of gravity, testing whether the GW luminosity distance is equal to the one of EM radiation has become the next goal. Current bright siren observations can only give weak constraints given the closeness of the sources [232], but spectral sirens can improve multi-messenger bounds given its larger redshift range [233]. There has been a lot of activity to constrain this modified propagation with different standard siren methods [234, 112, 235, 236, 237, 238, 239, 240] and to forecast their future capabilities with XG detectors [241, 242, 235, 243]. Besides these effects, future GW detectors will be able to constrain other phenomena. For example, they will bound the presence of additional cosmological tensor fields [244, 245], which have a rich phenomenology of waveform distortions [246], and test lensing effects beyond GR [247, 248]. Future GW observations combined with the redshifts from their host galaxies will also serve to measure peculiar velocities in order to constrain the growth of structure and test gravity [249].
@@ 5 Conclusions
Different standard siren methods will complement each other in the XG era, having capabilities that cannot be achieved with current detectors. On the one hand, very loud, nearby events will have such a good localization that their EM counterparts can be easily sought for or only one galaxy may lay within their sky map, granting a precise measurement of the local expansion rate. On the other hand, the large population of compact binaries will allow to narrowly constrain the expansion rate at high redshift. When put together, we project that standard siren cosmography has the potential to constrain the Hubble expansion rate $H(z)$ to sub-percent precision over more than 10 billion years of cosmic evolution, an accuracy that goes beyond planned cosmic surveys. This will be transcendental in multiple axes, having enough precision to arbitrate the tension among other cosmological measurements and to potentially discover new physics.
@@ Appendix A Detector locations and sensitivities
In sections 4.1 and 4.2, we consider five different detector networks and compare the bounds that can be put on $H_{0}$ with networks that contain these detectors. The names of these networks and the observatories that comprise them are given in table 2. The HLA network contains LIGO-H, LIGO-L and LIGO-A at ${}^{\sharp}$ sensitivities. We did not consider Virgo and KAGRA for this network as it is still unclear if these two detectors will upgrade to similar sensitivities to ${}^{\sharp}$, although their contributions at A+ level would still be very valuable for sky-localization estimates, aiding GW cosmology [102, 8]. ET contains the triangular ET configuration, while CE40ET considers a network with both the 40 km CE detector (at CE-A location) and the triangular ET observatory. We also consider a CE-only network that contains the 40 km CE (at CE-A) and the 20 km CE (at CE-B). Finally, the most advanced network contains the 40 km CE, 20 km CE and the triangular ET.
@@
HLA LIGO-L, LIGO-H, LIGO-A at ${}^{\sharp}$ sensitivity
@@ B.1 Mass and spin distribution
The population of BBHs is consistent with the inferred population with GWTC-3 events [163]. The population parameters for BNS and NSBH events are uncertain, owing to the lack of GW observations. The NS EOS is chosen to be APR4 [250]. Following the maximum mass limit applied by APR4, we choose a uniform mass distribution between $[1,2.2]M_{\odot}$ for neutron stars.
@@
The fiducial mass spectrum of BNSs and BBHs is presented in figure 10.
@@ B.2 Redshift distribution
with $\gamma=2.7$, $z_{p}=1.9$, and $\kappa=5.6$. While the most accurate approach would be to convolve a chosen time-delay distribution with the star formation rate to obtain the merger rate, we have decided against choosing a time-delay distribution for this work. This is because the time-delay distribution is still uncertain, and the Madau-Dickinson star formation rate by itself is consistent with the current bounds on the merger rate with GWTC-3 events [163]. The resultant redshift distribution is plotted in figure 11. Following the currently inferred rates [256, 163], the local merger rate densities for BBH, NSBH and BNS systems are chosen to be $\mathcal{R}_{0}^{\mathrm{BBH}}=24\,\,\mathrm{Gpc}^{-3}\,\mathrm{yr}^{-1}$, $\mathcal{R}_{0}^{\mathrm{NSBH}}=45\,\,\mathrm{Gpc}^{-3}\,\mathrm{yr}^{-1}$ and $\mathcal{R}_{0}^{\mathrm{BNS}}=320\,\,\mathrm{Gpc}^{-3}\,\mathrm{yr}^{-1}$, respectively.

### v2

@@ 4.2 Bright sirens
As seen in section 2.1, BNS mergers can lead to a variety of EM counterparts which can be detected by various EM observatories (see section 3.2). In this section, we focus on kilonovae that are detected by LSST for bright siren measurement. From the simulated population of detected BNS events, we choose the sub-population that lies within the redshift of $0.5$ and can be localized in the sky to $\Delta\Omega\leq 100\,\mathrm{deg}^{2}$, corresponding to each detector network. The choice of redshift corresponds to our expectation of the range of LSST for kilonova measurement, while the $\Delta\Omega$ cut-off ensures that LSST can observe the kilonova within $10$ sky patches. Then, using the kilonova model discussed in Refs. [229, 230, 231], we generate kilonova light curves in the $ugrizy$ bands of LSST, for all the events in these sub-populations. Assuming single-exposure with an exposure time of $30$ seconds, we claim that a kilonova is detected if its peak brightness exceeds the limiting magnitude of the telescope in that particular band. Using this approach, we find that most kilonova observations correspond to the $g-$band, with the maximum redshift at which a kilonova is detected to be $z=0.28$ (validating our $z\leq 0.5$ cut-off for the sub-populations). Further, we assume that only $40\%$ of these kilonovae will actually be detected [232]. This fraction aims to account for the LSST sky coverage, systematic biases as well as duty-cycles corresponding to the GW and EM observatories. The remaining set of events, those that are detected both in the GW and the EM band, are used for $H_{0}$ measurement using the same methodology that was used for golden dark sirens in section 4.1.
@@
We note that our bright siren estimates could be on the conservative side, as a result of our choice of the neutron star EOS, LSST detection strategy, or restricting ourselves to kilonova follow-up. In particular, we chose APR4 [233] to be the neutron star EOS, which results in relatively more compact neutron stars. Due to higher compactness, these neutron stars are difficult to disrupt tidally, reducing the amount of ejecta produced, which lowers the number of KN detections [118]). Further, several studies have underscored the significance of short gamma-ray burst detections in enhancing the precision of bright siren measurements [232, 234, 170, 235]. These studies particularly emphasize their efficacy in probing events at high redshifts, thereby contributing to constraints on $\Omega_{m}$. Moreover, while our focus has been on kilonovae observations BNS mergers, a valuable avenue lies in considering NSBH mergers, which can also produce kilonova counterparts, thereby extending the pool of potential bright sirens for $H_{0}$ measurement [131, 118]. Nevertheless, amidst these considerations, the overarching conclusion is that BNS bright sirens can, by themselves, achieve $H_{0}$ measurements with sub-percent accuracy. This underscores the potential of BNS bright sirens as a powerful tool for precise cosmological parameter estimation.
@@ 4.3 High-$z$ spectral sirens
We forecast the capabilities of spectral sirens performing a hierarchical Bayesian analysis on simulated data. We follow our fiducial population of BBHs and BNSs described in the appendices and simulate their detected posterior samples following the method described in [152]. We assume a flat $\Lambda$CDM cosmology whose expansion rate is described by $H_{0}$ and $\Omega_{m}$, cf equation (1) and 2, taking as fiducial values the ones reported in Planck 2018 [26]: $H_{0}=67.66$km/s/Mpc and $\Omega_{m}=0.31$. We take a 40km CE sensitivity as representative for XG detectors. The inference is performed using numpyro [236] and accelerated with JAX [237]. Our code is available on github. The inference is performed across all the cosmological and mass function parameters. We do not consider NSBHs as their mass spectrum is still largely unknown.
In figure 6 we present the projected constraints on the expansion rate of the Universe as a function of redshift. We compare the $1\sigma$ relative errors for a fixed number of BBHs and BNSs, taking the error with respect to the mean of the $H(z)$ traces at each redshift. Given current rate uncertainties, we choose a fixed number of 50,000 events which corresponds approximately to the detection rate of BBHs per year with XG detectors (see figure 3). Since we assumed the same merger rate history $\mathcal{R}(z)$ for all compact binaries (see details in B.2) and BNS selection effects brings them to lower redshifts, the population of BNSs has more constraining power at lower $z$. Within the population of BBHs, it is the peak at low masses that dominates the inference [152]. Both BNSs and BBHs can constrain the expansion rate with sub-percent precision over a wide range of redshifts, with the best constrained values at around $z\sim 1$. The best constrained redshift is located where most of the (well constrained) events are detected, which depends on the merger rate history and the detector’s sensitivity. A combined analysis of all CBCs will only improve these constraints and allow to more efficiently break the possible degeneracies with the redshift evolution of the mass spectrum [152].
@@ 4.4 All-in XG cosmography
Although we have introduced each type of standard siren separately for pedagogical reasons, they will all be analyzed simultaneously with XG detectors. In fact, as already discussed in the context of current detectors [136], the population inference of the spectral siren method must be intertwined with the dark siren method, since an incorrect population model assumption can bias the dark siren cosmology. Recently, there have been developments of algorithms to jointly infer the GW population and cosmology with additional galaxy catalog information using hierarchical Bayesian analyses [238, 239] and neural posterior estimation [240].
@@
Figure 7: Joint inference of spectral sirens and (golden) dark sirens with one year of observation of XG detectors. On the left, posterior distribution for $H_{0}$ and $\Omega_{m}$. The fiducial value is indicated with a dashed vertical line. On the right, $1\sigma$ relative errors in $H(z)$. The dark siren prior in $H_{0}$ has a $1\%$ error and we use BBHs as the reference population for spectral sirens.
@@ 4.5 Stress-testing the standard cosmological model
For the second class, many (exotic) variants are possible. One that can particularly be well tested with GWs is that general relativity (GR) does not hold at cosmological scales. Among other things, this changes the propagation of GWs. Since compact binary coalescence occur at cosmological distances, even very small deviations in the GW propagation with respect to GR can be strongly constrained. The paradigmatic modifications of gravity at cosmological scales that have been explored are those that change the speed of propagation and the scaling of the luminosity distance with redshift [243]. In particular, for the case of scalar-tensor theories, the tremendous power of tests of the speed of gravity were anticipated [244, 245] and spectacularly demonstrated with the first multi-messenger event, GW170817 [246, 247, 248, 249]. After the tight constraints on the speed of gravity, which cannot differ from the speed of light in more than 1 part in $10^{15}$, testing whether the GW luminosity distance is equal to the one of EM radiation has become the next goal. This modified luminosity distance can be parametrized in terms of an additional friction term in the propagation equation $\nu$, where GR predicts $\nu=0$. Current bright siren observations can only give weak constraints given the closeness of the sources [250], $|\nu|<\mathcal{O}(30)$, but spectral sirens can improve multi-messenger bounds given its larger redshift range [251], $|\nu|<\mathcal{O}(3)$. There has been a lot of activity to constrain this modified propagation with different standard siren methods [252, 133, 253, 254, 255, 256, 257, 258] and to forecast their future capabilities with XG detectors [259, 260, 253, 261], which could improve by another order of magnitude. Besides these effects, future GW detectors will be able to constrain other phenomena. For example, they will bound the presence of additional cosmological tensor fields [262, 263], which have a rich phenomenology of waveform distortions [264], and test lensing effects beyond GR [265, 266]. Future GW observations combined with the redshifts from their host galaxies will also serve to measure peculiar velocities in order to constrain the growth of structure and test gravity [267].
@@ 5 Conclusions
Different standard siren methods will complement each other in the XG era, having capabilities that cannot be achieved with current detectors. On the one hand, very loud, nearby events will have such a good localization that their EM counterparts can be easily sought for or only one galaxy may lay within their sky map, granting a precise measurement of the local expansion rate. On the other hand, the large population of compact binaries will allow to narrowly constrain the expansion rate at high redshift. When put together, we project that standard siren cosmography has the potential to constrain the Hubble expansion rate $H(z)$ to sub-percent precision over more than 10 billion years of cosmic evolution, a precision that is at least comparable to (or better than) the projections from other direct measurements [268] and some of the indirect measurements [269, 270, 271, 272]. This will be transcendental in multiple axes, having enough precision to arbitrate the tension among other cosmological measurements and to potentially discover new physics.
@@ Appendix A Detector locations and sensitivities
In sections 4.1 and 4.2, we consider five different detector networks and compare the bounds that can be put on $H_{0}$ with networks that contain these detectors. The names of these networks and the observatories that comprise them are given in table 2. The HLA network contains LIGO-H, LIGO-L and LIGO-A at A sensitivities. We did not consider Virgo and KAGRA for this network as it is still unclear if these two detectors will upgrade to similar sensitivities to A, although their contributions at A+ level would still be very valuable for sky-localization estimates, aiding GW cosmology [118, 7]. ET contains the triangular ET configuration, while CE40ET considers a network with both the 40 km CE detector (at CE-A location) and the triangular ET observatory. We also consider a CE-only network that contains the 40 km CE (at CE-A) and the 20 km CE (at CE-B). Finally, the most advanced network contains the 40 km CE, 20 km CE and the triangular ET.
@@
HLA LIGO-L, LIGO-H, LIGO-A at A sensitivity
@@ B.1 Mass and spin distribution
The population of BBHs is consistent with the inferred population with GWTC-3 events [185]. The population parameters for BNS and NSBH events are uncertain, owing to the lack of GW observations. The NS EOS is chosen to be APR4 [233]. Following the maximum mass limit applied by APR4, we choose a uniform mass distribution between $[1,2.2]M_{\odot}$ for neutron stars. The fiducial mass spectrum of BNSs and BBHs is presented in figure 10.
@@ B.2 Redshift distribution
with $\gamma=2.7$, $z_{p}=1.9$, and $\kappa=5.6$. While the most accurate approach would be to convolve a chosen time-delay distribution with the star formation rate to obtain the merger rate, we have decided against choosing a time-delay distribution for this work. This is because the time-delay distribution is still uncertain, and the Madau-Dickinson star formation rate by itself is consistent with the current bounds on the merger rate with GWTC-3 events [185]. Further, the redshift distribution is obtained by normalizing the star formation rate to match the local merger rate densities for the compact binaries. Following the currently inferred rates [278, 185], the local merger rate densities for BBH, NSBH and BNS systems are chosen to be $\mathcal{R}_{0}^{\mathrm{BBH}}=24\,\,\mathrm{Gpc}^{-3}\,\mathrm{yr}^{-1}$, $\mathcal{R}_{0}^{\mathrm{NSBH}}=45\,\,\mathrm{Gpc}^{-3}\,\mathrm{yr}^{-1}$ and $\mathcal{R}_{0}^{\mathrm{BNS}}=320\,\,\mathrm{Gpc}^{-3}\,\mathrm{yr}^{-1}$, respectively. The resultant redshift distribution is plotted in figure 11.